### v1

# Unified Thinker: A General Reasoning Modular Core for Image Generation

Abstract
Despite impressive progress in high-fidelity image synthesis, generative models still struggle with logic-intensive instruction following, exposing a persistent reasoning–execution gap. Meanwhile, closed-source systems (e.g., Nano Banana) have demonstrated strong reasoning-driven image generation, highlighting a substantial gap to current open-source models. We argue that closing this gap requires not merely better visual generators, but executable reasoning: decomposing high-level intents into grounded, verifiable plans that directly steer the generative process. To this end, we propose Unified Thinker, a task-agnostic reasoning architecture for general image generation, designed as a unified planning core that can plug into diverse generators and workflows. Unified Thinker decouples a dedicated Thinker from the image Generator, enabling modular upgrades of reasoning without retraining the entire generative model. We further introduce a two-stage training paradigm: we first build a structured planning interface for the Thinker, then apply reinforcement learning to ground its policy in pixel-level feedback, encouraging plans that optimize visual correctness over textual plausibility. Extensive experiments on text-to-image generation and image editing show that Unified Thinker substantially improves image reasoning and generation quality.
Unified Thinker: A General Reasoning Modular Core for Image Generation
[…]
Figure 1: Challenges in reasoning-aware image generation. Existing models, exemplified by Qwen-Image-Edit, exhibit two failure modes: (1) inaccurate reasoning (without Thinker), leading to logically incorrect edits; and (2) imprecise rendering (with Thinker), where correct reasoning does not translate into faithful visual outputs. Our Unified Thinker aims to address both issues.
Figure 2: Visual demonstrations of Unified Thinker on unified image generative tasks, including image editing and text-to-image generation, along with reasoning.
1 Introduction
The rapid evolution of diffusion-based foundation models ho2020denoising; dhariwal2021diffusion; rombach2022high; xu2024diffusion_dpo has driven an unprecedented leap in high-fidelity image synthesis. Advanced proprietary models such as GPT-4o hurst2024gpt4o and Nano Banana nanobanana have recently demonstrated strong reasoning-driven image generation under complex instructions. In contrast, despite steady progress in open-source systems sd3.5; blackforest2024flux; bagel2025; qwenimage-edit; liao2025imagegen-cot, current open-source models still exhibit a clear gap in handling logic-intensive or implicit directives wisebench; risebench; reason2edit; liu2025step1x-edit.
Current attempts to bridge this gap follow two primary approaches. Built-in Reasoning internalizes reasoning into the generator via unified training that couples multimodal understanding with generation bagel2025; show-o2; xiao2024omnigen. However, this tight entanglement reduces modularity and may destabilize training, often degrading the generator’s visual fidelity. In contrast, External Planner-Driven methods use an MLLM to plan for a mostly frozen generator qwenimage-edit; lin2025uniworld; li2025editthinker; reason2edit. While modular, they suffer from a reasoning–execution mismatch: text-space plans are not grounded in the generator’s capabilities, so even correct plans can cause visual failures, and iterative planning further increases compute.
We identify the key bottleneck as the absence of a principled paradigm for reasoning in image generation. In this paper, we propose Unified Thinker, a universal reasoning core for general image generation, built around a think-then-execute architecture that parametrically decouples the Thinker for instruction understanding and planning from the Generator for pixel synthesis. Here, the Generator refers to the underlying image synthesis backbone (e.g., a diffusion model) that takes conditioning signals and produces the final image in pixel space. The Thinker is implemented as a standalone, trainable multimodal large language model (MLLM) that transforms an instruction into a hierarchical, generator-friendly plan consisting of an intent summary, explicit constraints, and ordered sub-goals, which the Generator consumes as conditioning, enabling strong task transferability across text-to-image and editing and plug-and-play compatibility with different generator backbones.
However, this decoupled design still faces additional challenges: as shown in Fig. 1, without proper alignment, a naive Thinker may produce plausible reasoning that the Generator cannot execute. To bridge the reasoning-to-execution gap, we introduce a dedicated data-to-training pipeline to align planning with visual outcomes. We first construct HieraReason-40K, a hierarchical reasoning dataset synthesized with Gemini-3-Pro nanobanana, which pairs complex instructions with structured, executable plans to teach the Thinker the desired planning format and basic logical decomposition. We then adopt a two-stage training strategy: we perform joint supervised fine-tuning on HieraReason-40K to establish initial plan quality, followed by an end-to-end dual-phase reinforcement learning procedure that places the Generator in the loop and optimizes the Thinker using rewards computed from the final image’s constraint satisfaction. This directly grounds the Thinker’s policy in pixel-level feedback, encouraging plans that are not only semantically plausible but also executable under the Generator’s capabilities.
We conduct extensive evaluations in four settings: text-to-image reasoning, reasoning-based image editing, general text-to-image generation, and general image editing. Across all benchmarks, Unified Thinker delivers substantial gains in generative reasoning, markedly improving instruction following and constraint satisfaction. These improvements also hold across multiple generator backbones, supporting our core claim that a decoupled Thinker learns reusable, executable reasoning patterns that transfer across models and tasks.
Our main contributions are as follows:
  • We propose a decoupled reasoning-generation framework Unified Thinker that utilizes a unified module to handle general image generation tasks, significantly enhancing modular adaptability and transferability.
  • We introduce an end-to-end training pipeline spanning from hierarchical reason data construction to execution-led reinforcement learning, bridging the gap between abstract reasoning and pixel-level execution.
  • Through comprehensive experimental results, we demonstrated a significant performance improvement in reasoning-intensive generation tasks and verified the cross-model portability of our reasoning core module.
2 Related Work
2.1 Foundational Generative Models
Modern image generation is predominantly anchored in diffusion-based frameworks ho2020denoising; rombach2022high. Recent advances sd3.5; blackforest2024flux; qwenimage-edit build upon Diffusion Transformers peebles2023scalable and flow matching flowmatch to improve fidelity, prompt alignment, and diversity in latent diffusion models. Meanwhile, an emerging direction unifies autoregressive modeling with visual generation in a single framework, giving rise to unified multimodal models bagel2025; janus; show-o2; xiao2024omnigen. For instance, Bagel bagel2025 uses a transformer backbone to jointly model text and image tokens, whereas OmniGen xiao2024omnigen dispenses with external encoders and handles multiple vision tasks through a unified pipeline. In parallel, image editing has evolved from mask-based inpainting zhuang2024task; ju2024brushnet to instruction-guided manipulation brooks2023instructpix2pix; yu2025anyedit. To further enhance instruction following, recent methods li2024smartedit; fu2024guiding; lin2025uniworld; liu2025step1x-edit such as Qwen-Image-Edit qwenimage-edit leverage MLLMs for instruction parsing and planning. However, these models fall short in executing the complex logic required for sophisticated tasks, motivating us to introduce a dedicated Thinker module that bolsters the model’s fundamental reasoning capabilities during generation.
2.2 Reasoning for Image Generation
Recent research has moved beyond the one-shot mapping paradigm by explicitly incorporating reasoning into the image generation process. One line of work t2ir1; wang2025mint; liao2025imagegen-cot; unicot; huang2025interleaving introduces clear intermediate representations to decompose complex prompts into structured steps or explicit spatial layouts, improving compositional consistency and coherence. Another line of work reason2edit; mi2025thinkdiff; bagel2025 encourages models to reason about intent and constraints before drawing, moving beyond one-shot planning to better satisfy complex requirements, like R-Genie rgenie, which infers latent user intent instead of merely following the surface-level prompt. A third line of work guo2025canwegen; wu2025omnigen2; li2025reflectdit; li2025editthinker; yin2025reasonedit focuses on post-generation refinement by introducing reflection-and-correction mechanisms that assess the generated image, diagnose issues, and iteratively update the output to improve final quality. For example, Reflect-DiT li2025reflectdit introduces explicit self-reflection to guide revision, while EditThinker li2025editthinker enables reasoning via multi-round reflective interactions throughout the editing process. In contrast to these approaches, we propose a universal decoupled thinker that offers reusable reasoning as a standalone module, enabling easy transfer across diverse image generation and image editing tasks.
3 Data Construction
Goal and dataset.
We aim to train a standalone Thinker that augments existing diffusion generators with transferable reasoning while remaining generator-agnostic. To this end, we construct HieraReason-40K, a selected general-purpose corpus by combining four sources that cover text-to-image generation, general image editing, reasoning image generation, and reasoning image editing tasks han2025unireditbench; huang2025interleaving; qian2025pico; fang2025fluxreason. Each example pairs an instruction (optionally with reference images) with a structured reasoning trace that ends in an enhanced prompt for the downstream generator.
Figure 3: Data construction pipeline for HieraReason-40K. We combine seed knowledge and user requests to generate structured reasoning traces and executable enhanced prompts.
Structured reasoning trace.
As illustrated in Fig. 3, we create inference-style supervision by combining broad seed knowledge (e.g., art & culture) with input instruction to form generated inference data. Each training example is then rewritten into a rigorous structured reasoning trace: given an original instruction (and an optional reference image for image editing), the annotator produces a formalized reasoning trace followed by a final enhanced prompt for the generator. The trace follows a fixed three-stage procedure. First, it analyzes the input to identify the task type (text-to-image generation or image editing) and summarizes the intent. Next, it makes implicit requirements explicit and performs any necessary reasoning, such as counting, puzzle solving, numerical computation, temporal extrapolation, rule-based transformations, and attribute or coordinate lookup, to derive a concrete visual target. Finally, it converts the resolved target into an executable enhanced prompt. For image editing requests, we enforce an edit-only principle: the enhanced prompt describes only the intended changes, assuming all unspecified content is inherited from the reference image. This design ensures that reasoning is fully completed within the trace, while the downstream generator receives only a renderable visual specification.
Annotation and quality control.
We use Gemini3-Pro nanobanana to generate initial structured reasoning traces, followed by automatic normalization to enforce strict format consistency (e.g., mandatory stage headers and standardized image placeholders such as <image>). We further filter or rewrite samples that violate the trace format, fail to follow the edit-only principle for image editing, produce non-visual or underspecified targets, or exhibit inconsistencies between the reasoning trace and the final enhanced prompt. To further strengthen reasoning, we also carefully design a set of task-general system prompts that cover diverse common generation and editing scenarios.
4 Framework and Training
The core objective of our framework is to mitigate the reasoning–execution mismatch in reasoning-driven image generation and editing. We introduce a decoupled think-then-execute framework with two components: Thinker, a standalone, trainable multimodal large language model that produces structured reasoning traces and an executable visual specification, and Generator, a diffusion-based model that synthesizes the final image conditioned on the Thinker’s outputs. Training follows a two-stage pipeline, starting with joint supervised fine-tuning on structured traces and then moving to an execution-led, dual-phase reinforcement learning stage that optimizes the Thinker using rewards computed from the final generated images.
4.1 Joint Supervised Fine-Tuning
To teach the Thinker a consistent reasoning format and establish the think-then-execute pipeline, we first perform joint supervised fine-tuning stage. Given an instruction and an optional input image for editing, the Thinker produces a structured reasoning trace and an executable visual specification, and the Generator synthesizes the image conditioned on this output for both text-to-image generation and instruction-driven editing.
The training data is organized around instruction-following image generation and editing examples, each containing a user instruction, an optional reference image, and a target image. We derive two synchronized views of the same examples for joint training: (1) an understanding view, which pairs the input (instruction and optional reference image) with the annotated structured reasoning trace, supervising the Thinker via a language modeling loss; and (2) a generation view, which pairs the executable enhanced prompt (extracted from the trace) with the target image, supervising the Generator via the standard diffusion denoising objective. During each training step, we sample mini-batches from the two views and optimize a weighted sum of the understanding loss $\mathcal{L}_{\text{und}}$ (token-level cross-entropy) and the generation loss $\mathcal{L}_{\text{gen}}$ (noise-prediction mean squared error).
Figure 4: Our proposed two-stage framework for reasoning-aware image generation. Stage 1 initializes the Thinker Model and Generator Model. Given an Image & Prompt (x), the Thinker generates a Reasoning Thought (y), which then guides the Generator to produce a Refined Image (z). Stage 2 further refines the Thinker and Generator Models to enhance their capability in integrating complex reasoning (y) into high-fidelity visual outputs (z), applicable to both novel image generation and existing image editing tasks.
This joint supervised fine-tuning procedure effectively aligns the instruction-generation capability of the Thinker with the image-synthesis prior of the Generator, ensuring that the produced reasoning instructions are not only semantically accurate but also highly compatible with the Generator’s operational semantics, thereby laying a solid foundation for cascaded inference deployment.
Formally, the overall objective is defined as:
$\mathcal{L}_{\text{SFT}}=$ $\mathcal{L}_{\text{gen}}\big(\text{Generator}(\mathbf{y},\,\mathbf{x}_{\text{ref}}),\,\mathbf{x}_{\text{tgt}}\big)$ (1)
$+\lambda\,\mathcal{L}_{\text{und}}\big(\text{Thinker}(\mathbf{x}_{\text{img}}),\,\mathbf{y}\big),$
where $\mathbf{x}_{\text{img}}$ denotes the input image, $\mathbf{y}$ is the ground-truth reasoning process, $\mathbf{x}_{\text{ref}}$ represents an optional reference image, $\mathbf{x}_{\text{tgt}}$ is the target output image, and $\lambda>0$ is a hyperparameter balancing the two learning signals.
4.2 Dual-Phase Reinforcement Learning
While joint fine-tuning provides an initial alignment, it leaves a nontrivial reasoning–execution gap: the Thinker may produce plans that are plausible in text but suboptimal for the generator to execute. To address this without additional manual annotation, we introduce a dual-phase reinforcement learning strategy based on Group Relative Policy Optimization guo2025deepseekr1. The key idea is to sample multiple candidate traces for the same request, execute them with the generator, and train the Thinker by relative advantage feedback, promoting outputs that lead to better images and suppressing those that do not.
Phase 1: Reasoning-Oriented RL. In this phase, we optimize the Thinker’s ability to provide effective guidance. For a given instruction, the Thinker samples a group of $G$ reasoning paths $\{y_{1},y_{2},\dots,y_{G}\}$. We use the Generator (fixed) to produce the corresponding images and assign a reward $r_{i}$ to each path based on the final image quality. We optimize the Thinker by maximizing:
$\mathcal{J}_{T}(\theta_{T})=\mathbb{E}\left[\frac{1}{G}\sum_{i=1}^{G}\left(\frac{\pi_{\theta}(y_{i}|p)}{\pi_{old}(y_{i}|p)}\cdot\hat{A}_{i}\right)\right]$ (2)
$\hat{A}_{i}=\frac{r_{i}-\text{mean}(\{r\})}{\text{std}(\{r\})}$ (3)
Here, $\hat{A}_{i}$ is the relative advantage, which tells the model which reasoning chains performed better than the group average. This forces the Thinker to prioritize logic that is not just "correct" in text, but "useful" for the Generator.
Phase 2: Generation-Oriented RL. With the Thinker providing reliable plans, we next improve the Generator’s execution fidelity. However, probability-flow ODE sampling in diffusion models is essentially deterministic, limiting the stochastic rollouts required by reinforcement learning. Following a Flow-GRPO-like idea flowgrpo, we convert the ODE sampler into an equivalent reverse-time SDE to introduce controlled randomness, enabling $G$ distinct rollouts $\{z_{1},z_{2},\dots,z_{G}\}$ for the same instruction and optimizing the Generator accordingly:
$\mathcal{J}_{G}(\theta_{G})=\mathbb{E}\left[\frac{1}{G}\sum_{i=1}^{G}\left(\frac{\pi_{\theta}(z_{i}|c,p)}{\pi_{old}(z_{i}|c,p)}\cdot\hat{A}_{i}\right)\right]$ (4)
In this stage, the advantage $\hat{A}_{i}$ assigns higher credit to denoising trajectories that yield better images. With this two-stage feedback, the Thinker improves planning while the Generator improves execution, leading to substantial performance gains. Details of the reward design are provided in the appendix A.4.
Table 1: Performance comparison of models on the RiseBench benchmark. We report three general performance metrics: Instruction Reasoning (Reason.), Appearance Consistency (Consist.), and Visual Plausibility (Visual.). Additionally, we present category-wise accuracy (%) for four specific reasoning dimensions: Temporal, Causal, Spatial, and Logical. The Overall score is the average of these four category-wise accuracies.
Model Reason. Consist. Visual. Temporal Causal Spatial Logical Overall
Gemini-3-pro-image-preview 77.0 85.5 94.4 41.2 61.1 48.0 37.6 47.2
Gemini-2.5-Flash-Image 61.2 86.0 91.3 25.9 47.8 37.0 18.8 32.8
GPT-Image-1 62.8 80.2 94.9 34.1 32.2 37.0 10.6 28.9
GPT-Image-1-mini 54.1 71.5 93.7 24.7 28.9 33.0 9.4 24.4
Gemini-2.0-Flash-exp 48.9 68.2 82.7 8.2 15.5 23.0 4.7 13.3
BAGEL (w/ CoT) 45.9 73.8 80.1 5.9 17.8 21.0 1.2 11.9
Seedream-4.0 58.9 67.4 91.2 12.9 12.2 11.0 7.1 10.8
Gemini-2.0-Flash-pre 49.9 68.4 84.9 10.6 13.3 11.0 2.3 9.4
FLUX.1-Kontext-Dev 26.0 71.6 85.2 2.3 5.5 13.0 1.2 5.8
Ovis-U1 33.9 52.7 72.9 1.2 3.3 4.0 2.4 2.8
Step1X-Edit 30.3 12.6 74.9 0.0 2.2 2.0 3.5 1.9
OmniGen 25.1 41.5 73.5 1.2 1.0 0.0 1.2 0.8
EMU2 22.6 38.2 78.3 1.2 1.1 0.0 0.0 0.5
BAGEL 36.5 53.5 73.0 2.4 5.6 14.0 1.2 6.1
+ Unified Thinker (Qwen2.5-VL-7B) 53.3 73.6 78.1 14.1 17.7 18.0 3.5 13.6
+ Unified Thinker (Qwen3-VL-8B) 58.7 75.7 80.9 15.2 17.7 20.0 8.2 15.5
Qwen-Image-Edit 37.2 66.4 86.9 4.7 10.0 17.0 2.4 8.9
+ Unified Thinker (Qwen2.5-VL-7B) 58.6 75.9 90.1 24.7 22.2 38.0 9.4 24.2
+ Unified Thinker (Qwen3-VL-8B) 61.9 76.2 90.5 32.9 30.0 41.0 9.4 28.9
Table 2: Results on GEditBench for general instruction-based image editing. We report G_SC, G_PQ, and G_O on the English split.
Model G_SC $\uparrow$ G_PQ $\uparrow$ G_O $\uparrow$
UniWorld-V2 8.29 8.02 7.83
Step1x-edit-v1p2(reflection) 8.18 7.85 7.58
Step1x-edit-v1p2(thinking) 8.02 7.64 7.36
Step1X-edit-v1.1 7.66 7.35 6.97
Flux-Kontext-dev 7.16 7.37 6.51
OmniGen2 7.16 6.77 6.41
OmniGen 5.96 5.89 5.06
AnyEdit 3.18 5.82 3.21
BAGEL 7.36 6.83 6.52
+ Unified Thinker (Qwen2.5-VL-7B) 7.29 6.88 6.53
+ Unified Thinker (Qwen3-VL-8B) 7.38 6.75 6.60
Qwen-Image-Edit 8.00 7.86 7.56
+ Unified Thinker (Qwen2.5-VL-7B) 8.17 7.94 7.67
+ Unified Thinker (Qwen3-VL-8B) 8.15 8.04 7.71
Table 3: Results on PRISM for general text-to-image generation. We report alignment (Aln), aesthetics (Aes), and average (Avg) using GPT-4.1 as evaluation.
Model Aln $\uparrow$ Aes $\uparrow$ Avg $\uparrow$
Gemini-2.5-Flash-Image 87.1 83.4 85.3
Qwen-Image 81.1 78.6 79.9
SEEDream 3.0 80.5 78.7 79.6
HiDream-I1-Full 76.1 75.6 75.9
FLUX.1-Krea-dev 74.3 75.1 74.7
SD3.5-Large 73.9 73.5 73.7
FLUX.1-dev 72.4 74.9 73.7
HiDream-I1-Dev 70.3 70.0 70.2
BAGEL 66.7 63.4 65.1
+ Unified Thinker (Qwen2.5-VL-7B) 73.5 67.7 70.6
+ Unified Thinker (Qwen3-VL-8B) 75.1 69.2 72.1
Qwen-Image-Edit 76.9 70.7 73.8
+ Unified Thinker (Qwen2.5-VL-7B) 77.3 73.8 75.6
+ Unified Thinker (Qwen3-VL-8B) 83.2 73.0 78.1
Table 4: Results on WiseBench for reasoning-based text-to-image generation. We report accuracy across six knowledge domains and the overall score.
Model Cultural Time Space Biology Physics Chemistry Overall
GPT-4o 0.81 0.71 0.89 0.83 0.79 0.74 0.80
Qwen-Image 0.62 0.63 0.77 0.57 0.75 0.40 0.62
UniWorld-V2 0.60 0.61 0.70 0.53 0.64 0.32 0.58
UniWorld-V1 0.53 0.55 0.73 0.45 0.59 0.41 0.55
Manzano-3B 0.42 0.51 0.59 0.45 0.51 0.32 0.46
Manzano-30B 0.58 0.50 0.65 0.50 0.55 0.32 0.54
OpenUni-B-512 0.37 0.45 0.58 0.39 0.50 0.30 0.43
OpenUni-L-512 0.51 0.49 0.64 0.48 0.63 0.35 0.52
OpenUni-L-1024 0.49 0.53 0.69 0.49 0.56 0.39 0.52
MetaQuery-XL 0.56 0.55 0.62 0.49 0.63 0.41 0.55
Liquid 0.38 0.42 0.53 0.36 0.47 0.30 0.41
BAGEL 0.44 0.55 0.68 0.44 0.60 0.39 0.52
+ Unified Thinker (Qwen2.5-VL-7B) 0.72 0.65 0.75 0.64 0.75 0.61 0.70
+ Unified Thinker (Qwen3-VL-8B) 0.70 0.65 0.73 0.62 0.73 0.55 0.68
Qwen-Image-Edit 0.62 0.63 0.77 0.57 0.75 0.40 0.62
+ Unified Thinker (Qwen2.5-VL-7B) 0.75 0.66 0.78 0.75 0.79 0.61 0.73
+ Unified Thinker (Qwen3-VL-8B) 0.75 0.70 0.81 0.73 0.81 0.55 0.74
5 Experiments
We evaluate Unified Thinker in four settings: general instruction-driven image editing, general text-to-image generation, reasoning-intensive image editing, and reasoning-intensive text-to-image generation. Our goal is to examine whether the decoupled Thinker-Generator architecture, further strengthened by our dual-phase reinforcement learning, yields consistent gains over strong open-source baselines in instruction following and reasoning-grounded visual synthesis.
5.1 Experimental Setup
Model configuration. Unless otherwise specified, Unified Thinker uses Qwen2.5-VL-7B, and we additionally report results with Qwen3-VL-8B bai2024qwenvl; bai2025qwen3vltechnicalreport.We use Qwen-Image-Edit qwenimage-edit as the base generator to execute the visual specifications produced by the Thinker. For reinforcement learning and automated evaluation, we adopt Qwen3-VL-30B bai2025qwen3vltechnicalreport as the reward model, which provides feedback on both visual correctness and logical consistency.
Training data and setup. For the supervised cold start, we jointly fine-tune on HieraReason-40K. For reinforcement learning, we sample 4K high-quality instances from HieraReason-40K and apply Group Relative Policy Optimization (GRPO) to improve the Thinker’s structured outputs and their executability, thereby strengthening the Generator’s adherence to the resulting specification. Training uses NVIDIA H20 GPUs, with 16 GPUs for supervised fine-tuning and 64 GPUs for reinforcement learning.
Evaluation benchmarks. We evaluate on WiseBench wisebench, RISEBench risebench, GEditBench liu2025step1x-edit, and PRISMBench fang2025fluxreason. These benchmarks cover diverse knowledge domains and editing operations, requiring models to combine high-level semantic reasoning (e.g., temporal, and logical inference) with low-level visual manipulation (e.g., content preservation).
5.2 Main Results
Reasoning-based image editing (RISEBench). As shown in Table 1, our method markedly improves reasoning-heavy editing over the base Qwen-Image-Edit and a naive MLLM-thinker baseline. In particular, the unified training strategy yields large improvements on temporal and spatial reasoning, indicating that the Thinker effectively resolves hidden constraints (e.g., temporal shifts or relational edits) and reduces semantic drift during diffusion execution.
Text-to-image reasoning (WiseBench). Table 4 shows that Unified Thinker achieves the strongest overall performance among open-source models and improves most domain categories, substantially narrowing the gap to closed-source frontier models such as GPT-4o. Gains are especially notable in categories that demand precise entity grounding and knowledge retrieval (e.g., cultural and biology), suggesting that explicit planning helps translate implicit constraints into executable visual specifications.
General generation and editing quality. Beyond reasoning-centric benchmarks, we further confirm that incorporating the Thinker does not compromise general-purpose generation or editing performance. On PRISM (Table 3), our method achieves a consistent improvement in overall quality, with gains that are mainly reflected in aesthetic preference while preserving prompt-image alignment. On GEditBench (Table 2), Unified Thinker also delivers modest yet consistent gains across all reported metrics. Together, these results suggest that the planning stage improves instruction decomposition and visual target specification without weakening the Generator’s core rendering ability, and can even provide small benefits under standard, non-reasoning workloads.
5.3 Ablation Study
Training stage ablation. We conduct ablation studies on RiseBench, WiseBench, and GEdit, using Qwen-Image-Edit as the baseline and progressively adding the Thinker module, joint fine-tuning, and two-stage Dual-RL training.
Table 5 shows that introducing the Thinker notably improves performance on reasoning-oriented benchmarks, while slightly hurting low-level editing quality on GEdit, revealing a mild objective trade-off. Joint fine-tuning alleviates this mismatch and stabilizes multi-task behavior, and the proposed two-stage Dual-RL further yields consistent gains across all benchmarks, leading to the best overall results by better aligning reasoning with final visual outcomes.
Thinker backbone ablation. We instantiate Unified Thinker with two backbones(Qwen2.5-VL-7B and Qwen3-VL-8B). Overall, a stronger Thinker backbone tends to yield better reasoning-oriented performance and improves overall editing fidelity on reasoning tasks, whereas the 7B variant can be slightly preferred on PRISM in terms of aesthetics, suggesting a trade-off between logic alignment and visual preference. Moreover, Table 6 shows that using an external Thinker (regardless of the specific backbone) consistently outperforms the Qwen-Image-Edit baseline, with most gains coming from improved reasoning and consistency while visual quality remains comparable.
Table 5: Training stage ablation results on RiseBench, WiseBench, and GEdit. The baseline is based on Qwen-Image-Edit. The Thinker is implemented with Qwen2.5-VL-7B and further trained in our framework.
Ablation Rise $\uparrow$ Wise $\uparrow$ GEdit $\uparrow$
baseline 8.9 0.62 7.56
+ Thinker 16.4 0.66 7.49
+ Joint fine-tune 20.2 0.68 7.52
+Dual-RL stage 1 21.9 0.72 7.61
+ Dual-RL stage 2 24.2 0.73 7.67
Table 6: Ablation of the Thinker design on RiseBench. We report Reason., Consist., Visual., and Overall, where Overall is the average accuracy over Temporal, Causal, Spatial, and Logical. The baseline is Qwen-Image-Edit.
Model Reason. Consist. Visual. Overall
baseline 37.2 66.4 86.9 8.9
+ Gemini-2.5-Pro 64.3 71.9 88.3 25.2
+ GPT-5 67.4 76.6 86.3 26.9
+ Qwen3-VL-30B 57.6 75.9 86.6 23.1
+ Unified Thinker (7B) 58.6 75.9 90.1 24.2
Cross-generator ablation. We evaluate the transferability of the Thinker module by applying the Unified Thinker trained with the Qwen-Image-Edit pipeline to a different generator (BAGEL). As shown in Tables 1, 4, and 2, adding Thinker consistently improves BAGEL on both RiseBench and GEditBench, demonstrating that the module generalizes beyond the training generator and can be integrated into other generation models with stable gains.
6 Conclusion
We propose Unified Thinker, a decoupled Thinker–Generator framework that equips diffusion models with transferable reasoning and planning. The Thinker maps user requests for both text-to-image generation and image editing into a structured, executable intermediate representation, enabling the Generator to focus on faithful visual synthesis. We build a $\sim$40K cold-start training corpus with strict formatting and further enhance both planning and execution with a two-stage RL pipeline. Extensive experiments show consistent gains over strong open-source baselines, especially on reasoning-intensive requests, demonstrating the value of separating reasoning from rendering.
7 Limitations
Our approach still depends on the quality and coverage of the intermediate representation, training data, and automatic rewards used during RL, which can introduce bias and limit generalization beyond the evaluated benchmarks. While the Thinker is designed to be generator-agnostic, executability is not fully invariant across different diffusion backends, and some difficult edits (e.g., fine-grained geometric changes, strict locality, or precise text rendering) remain challenging. Finally, the additional planning stage increases inference latency and compute cost compared to directly prompting a single generator.
Appendix A Implementation Details
A.1 Joint Supervised Fine-Tuning
We perform joint supervised fine-tuning on Qwen2.5-VL-7B-Instruct and Qwen3-VL-8B-Instruct with LoRA (r=8r{=}8, applied to all modules) on 16 NVIDIA H20 GPUs,. Training uses a mixed instruction dataset (mixed edit and text-to-image data in HieraReason-40K) with the qwen3_vl template, maximum sequence length 8096. We use batch size 4 per device with 8 gradient accumulation steps, learning rate $4\times 10^{-5}$, cosine schedule with 10% warmup, for 5 epochs. We set $\lambda=0.5$. The image is resized so that the short side is 512 pixels, with aspect ratio preserved.
A.2 Dual-Phase Reinforcement Learning
We further optimize the models with GRPO on 64 GPUs. For rollouts, we use a batch size of 16 and generate 24 candidates per prompt with sequence expansion enabled. We sample outputs with top k=100, and temperature 0.99, allowing up to 8192 new tokens; both prompt and response are capped at 8192 tokens. Each iteration performs one update epoch with clipping thresholds of 0.5 (value), 10 (reward), and 10 (advantage), without advantage whitening. We include KL regularization with a coefficient of 0.01 against a reference model. The ’thinker’ actor is initialized from Qwen2.5-VL-7B-Instruct/Qwen3-VL-8B-Instruct and trained in BF16 using a learning rate of $1\times 10^{-6}$ and weight decay 0.01, with an effective batch size realized via 1 sample per GPU and 96 gradient accumulation steps under Megatron parallelism (tensor parallelism 4 with sequence parallelism). Rewards are computed using Qwen3-VL-30B-A3B-Instruct as the VLM judge and Qwen-Image-Edit as the editor, using 10 edit sampling steps.
A.3 Training Evaluation
Fig. 5 plots the mean reward score during training, which increases steadily, indicating consistent improvement of the learned policy. Fig. 6 reports the per-step rollout generation time, showing the runtime behavior throughout training.
A.4 Reward Model and Design
We use a VLM-based reward model to provide a scalar supervision signal for both image-editing and text-to-image (T2I) training. For image editing, the judge is conditioned on the pre-edit image, the post-edit image, the edit instruction (edit_prompt), and a reference description of the intended outcome (edit_prompt_cot), and returns three integer subscores on a 1–5 scale: Appearance Consistency (whether non-instructed regions remain unchanged), Reasoning/Alignment (how well the edited image matches the intended result under the instruction), and Visual Plausibility (realism and overall generation quality); these subscores are aggregated into a single scalar reward. For T2I, we first synthesize an image from the prompt (and, when applicable, the answer field extracted from the model output) and then evaluate it with the same VLM judge using a strict rubric that outputs three integer subscores in $\{0,1,2\}$—Consistency (prompt-image alignment), Realism (physical plausibility and fidelity), and Aesthetic Quality (overall visual appeal)—whose mean yields the final reward in $[0,2]$ for reinforcement learning.
Figure 5: Mean reward score over training.
Figure 6: Per-step rollout generation time over training.
Source Samples Task Type Input Output
Unireditbench $\sim$ 10K Reasoning Image editing instruction + image think + enhanced prompt
Pico-Banana-400K $\sim$ 10K Image editing instruction + image think + enhanced prompt
IRGL-300K $\sim$ 10K Reasoning T2I instruction think + enhanced prompt
Flux-reason-6m $\sim$ 10K T2I instruction think + enhanced prompt
Total 40K 4 categories think + enhanced prompt
Table 7: Composition of HieraReason-40K. We sample 10K instances from each source dataset and distill them into a unified, structured format using Gemini with our system prompt.
A.5 Evaluated Comparative Models
The following models were used in our comparative evaluation:
Gemini-2.5-Flash-Image(Nano Banana): A state-of-the-art multimodal model by Google optimized for high-fidelity text-to-image generation, complex image editing, and multi-image composition nanobanana.
GPT-Image-1 & GPT-4o: OpenAI’s unified multimodal series that demonstrates advanced spatial-temporal reasoning and end-to-end processing across text and vision hurst2024gpt4o.
Gemini-2.0-Flash: A multimodal model from Google designed for real-time visual and textual reasoning tasks geminifamily.
BAGEL: A unified understanding and generation multimodal framework that incorporates Chain-of-Thought (CoT) reasoning to improve logical deduction in visual tasks bagel.
Qwen-Image / Edit: A series of vision-language models from Alibaba; the Edit variant is specifically fine-tuned for instruction-based image manipulation qwenimage-edit.
EMU2: A generative multimodal model that uses a unified modeling framework for both visual-sequential understanding and generation emu2.
FLUX.1 (Dev/Kontext): A flow-matching based rectified flow transformer model known for superior text rendering and adherence to complex prompts blackforest2024flux.
Stable Diffusion 3.5 (SD3.5): A Multimodal Diffusion Transformer (MMDiT) architecture optimized for high-resolution synthesis and prompt following sd3.5.
OmniGen: A unified image generation model capable of handling various tasks including generation, editing, and control within a single framework wu2025omnigen2.
Ovis: An open-source structural visual-language model designed to process high-resolution images with structural integrity ovis.
Step1X-Edit (v1.1/v1.2): A family of generative models by StepFun; the v1.2 variants utilize "thinking" and "reflection" mechanisms to improve reasoning-heavy editing tasks liu2025step1x-edit.
UniWorld (V1/V2): A multimodal world model framework designed for spatial-temporal understanding and high-fidelity video/image synthesis lin2025uniworld.
Manzano : A unified multimodal large model framework with a shared visual encoder.MANZANO.
OpenUni (B/L): A fully open-source lightweight multimodal unified baseline. It connects existing multimodal large language models with diffusion models through learnable queries and a lightweight Transformer connector, thereby enabling simultaneous multimodal understanding and image generation. openuni.
MetaQuery-XL: An expanded multimodal. It connects the frozen multimodal large model and the diffusion model with a set of learnable queries, transferring the understanding and reasoning capabilities of the large model to image generation. metaqueriesxl.
Liquid: An extensible unified autoregressive generation paradigm that discretizes images into tokens and shares the same token/embedding space with text tokens, enabling a single large language model to simultaneously perform multimodal understanding and image generation. liquid.
Appendix B System Prompt
Table 8:
System Prompt  You are a Visual-Language Model (VLM) Prompt Optimization Expert specializing in image generation and editing. Your core task is to receive user instructions (potentially including a reference image), and after deep visual analysis and logical reasoning, output an enhanced English prompt (enhanced_prompt) for downstream Diffusion Models to generate high-quality images. ### Three Core Principles (Guiding Principles) You must always adhere to the following three unshakeable principles, which are the foundation of all your actions. 1. Task Dichotomy: Your primary judgment is to distinguish between "Text-to-Image (T2I)" and "Image-to-Image (I2I)."
- T2I is fundamentally about Creation: Your ‘answer‘ must describe the entire scene in detail from scratch.
- I2I is fundamentally about Modification: Your ‘answer‘ must be a precise instruction, describing only the change that needs to occur.
2. The "Golden Rule" for I2I (Modification Focus Principle): For any I2I task, your ‘answer‘ is strictly forbidden from containing descriptions of any areas or elements that should remain unchanged. The downstream model relies on the reference image to maintain constancy; restating these elements in the prompt will only lead to confusion and inconsistency.
3. The "Brain vs. Hand" Principle for Reasoning: If the task requires logical reasoning, calculation, knowledge retrieval, or conceptual transformation, you must act as the "Brain."
- Complete all thinking within the ‘<think>‘ tag and arrive at a concrete, visual final result.
- In the ‘<answer>‘ tag, you must directly provide the visual description of this result, rather than asking the "Hand" (the downstream Diffusion Model) to repeat your thinking process.
### Guide for Thinking Process (<think> Tag Content) You must structure your thinking within the ‘<think>‘ tag by naturally deconstructing the task through answering the following series of questions: Step 1: Input Analysis & Intent Identification Basic Judgment: Is this task "Text-to-Image" or "Image-to-Image"?
Intent Verb: What is the user’s core intent? Is it Add, Change, Replace, Isolate/Extract, Combine, Transform (style/pose/concept), or Solve/Draw (solve and then draw)?
Step 2: Reasoning Activation & Result Concretization Reasoning Check: Does fulfilling the intent from the previous step require reasoning beyond the literal meaning?
Execute Reasoning (If required): Immediately perform the required reasoning here.
Result Statement: After reasoning is complete, you must explicitly state: "The concrete visual result of my reasoning is: [Write the specific, visual answer here]".
Step 3: Strategy Formulation & Prompt Construction Comprehensive Decision: Formulate the final ‘answer‘ based on the "Task Type" (T2I/I2I), the "User Intent Verb," and the "Concrete Reasoning Result" (if any).
Principle-Based Construction:
- If the task is "Text-to-Image": enrich the scene from scratch.
- If the task is "Image-to-Image": describe only the change; refer to the given image.
### Output Format (<answer> Tag Content) Directly output a block of text, which must strictly adhere to the following format: <think> ... </think> <answer>Enhanced English Prompt</answer>
We design a system prompt that converts user instructions (optionally with a reference image) into high-quality English prompts for diffusion models. It enforces a strict T2I/I2I split: T2I describes the full scene, while I2I specifies only the required edits. A “golden rule” forbids restating unchanged content to reduce edit drift. Moreover, the “Brain vs. Hand” principle confines reasoning to <think> and outputs only the concrete visual result in <answer>.
This design supports four common scenarios: (1) T2I generation with complete scene specification; (2) I2I local edits (add/change/replace) with improved consistency; (3) combine/transform tasks via consolidated, non-conflicting visual descriptions; and (4) solve/draw tasks by forcing reasoning to be resolved into an explicit visual target before generation.
Appendix C Details of HieraReason-40K
HieraReason-40K is built to train a generator-agnostic Thinker that produces structured reasoning traces and a final enhanced prompt for downstream diffusion models. We collect 40K instruction examples from four open-source datasets covering image editing and reasoning-oriented generation/editing tasks han2025unireditbench; huang2025interleaving; qian2025pico; fang2025fluxreason. Specifically, we sample 10K instances from each source, and then convert them into a unified format via structured knowledge distillation with Gemini under our system prompt, yielding intermediate reasoning traces aligned with the final enhanced prompt.
Figure 7: Visual demonstrations of UnifiedThinker on unified image generative tasks.
Figure 8: Visual demonstrations of UnifiedThinker on unified image generative tasks.
Figure 9: Visual demonstrations of UnifiedThinker on unified image generative tasks.